Write a quadratic equation in the variable x having the given numbers as solutions.Type the equation in standard form.
The solutions are -3,1

Answer by ewatrrr(24785)

standard form: ax^2 + bx + c = 0
y = (x+3)(x-1)
(x+3)(x-1) = 0
x^2 + 2x - 3 = 0
F First terms
O Outside terms -x
I Inside terms 3x
L Last terms